OpenStudy (anonymous):

HELP!!! Consider the equations I. v = v0 + ax II. y = (2m) cos(kx), where k = 2m^-1 III. v^2 = ax Which ones are dimensionally correct?

OpenStudy (theeric):

Well, I haven't heard of "dimensionally correct," but you can check the units to find which ones \(can't\) be true!

OpenStudy (anonymous):

How? I am so confused with this problem!

OpenStudy (anonymous):

This are dimensions: Distance (x) = [L] Area = [L^2] Volume = [L^3] Velocity (v) = [L]/[T] Acceleration (a) = [L]/[T^2] Energy = [M][L^2]/[T^2]

OpenStudy (anonymous):

These are the dimensions:*

OpenStudy (theeric):

It is confusing, because I don't know what this is! Sorry!

OpenStudy (theeric):

I'm researching now.

OpenStudy (anonymous):

Thank you!

OpenStudy (shane_b):

Dimensionally correct is just saying that the units work out correctly.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I think we should skip that, since I still have two more questions, which I think are easier to deal with.

OpenStudy (anonymous):

While John is traveling along a straight interstate highway, he notices that the mile marker reads 257 km. John travels until he reaches the 149 km marker and then retraces his path to the 177 km marker. What is John's resultant displacement from the 257 km marker? Answer in units of km.

OpenStudy (shane_b):

Before you move on, just plug in the dimensions for each of the 3 equations (ie: v = L/T so put L/T in place of v...and so on). Then simplify the results and see which one is a valid equation...being that both sides are equivalent.

OpenStudy (shane_b):

For John's trip, draw it out first:

OpenStudy (anonymous):

That's what I've been doing, and I know that the first one isn't correct, but I don't know about the second and third one.

OpenStudy (shane_b):

So what's his total displacement?

OpenStudy (anonymous):

Should I subtract 257 - 177 to find his resultant displacement?

OpenStudy (shane_b):

Exactly.

OpenStudy (anonymous):

So, will his resultant displacement from the 257 km marker will be 80 km?

OpenStudy (anonymous):

So, will his resultant displacement from the 257 km marker be 80 km?*

OpenStudy (shane_b):

Yes. Going back to the other one:\[v^2 = ax\]Plug in values to get:\[(\frac{L}{T})^2=\frac{L}{T^2}*L\]So that's equivalent.

OpenStudy (shane_b):

Just do the same thing with the other 2 equations and see how it turns out.

OpenStudy (anonymous):

So, the third one is dimensionally correct, right?

OpenStudy (shane_b):

Yes...and the first one is not equivalent. Now try the second one

OpenStudy (anonymous):

I'm pretty sure the second one is dimensionally correct.

OpenStudy (shane_b):

How did you come to that conclusion?

OpenStudy (shane_b):

Before you can answer that you need to know how y is defined. Do you have that info?

OpenStudy (anonymous):

I actually did some research over the internet and found is dimensionally correct.

OpenStudy (anonymous):

I'm very confused with the second one, since it is y = (2m) cos(kx), where k=2m^-1.

OpenStudy (shane_b):

But you can figure this out on your own....it's very easy but I don't know what y is supposed to represent since it's not in your Dimensions post.

OpenStudy (anonymous):

That's why I'm confused.

OpenStudy (shane_b):

Here's a hint, the cosine of any value will be in what range?

OpenStudy (anonymous):

-1 to 1.

OpenStudy (shane_b):

So no matter what is inside those parenthesis, the result will be from 1 to -1...with no units. Therefore, you can actually forget about the cosine part of the equation altogether. Then you're left with this:\[y = 2m\]If y's units are the same as m's units, then the equation is valid.

OpenStudy (anonymous):

But, what are the units for y

OpenStudy (anonymous):

?

OpenStudy (shane_b):

I have no idea...it's not in the question. If you found it on the internet, maybe that question actually has more info in it.

OpenStudy (shane_b):

If y is something like the typical vertical distance (meters), then the equation would be considered dimensionally correct since the units on both sides are equal.

OpenStudy (shane_b):

That's probably a safe assumption in this case but I'd argue that the question is missing info.

OpenStudy (anonymous):

Thank you! Now, going back to the displacement question, I'm not sure about the answer I got.

OpenStudy (shane_b):

What isn't clear to you in that one? Think about the definition of displacement.

OpenStudy (anonymous):

I just find it too simple to be truth.

OpenStudy (anonymous):

I mean, 257-177 = 80, and 80 is the final answer?

OpenStudy (shane_b):

The question is intended to test your understanding of displacement versus distance...that's all. Yes, the final answer is 80km.

OpenStudy (anonymous):

But, what happened with the 149 km he reaches before he retraced his displacement to the 177 km marker.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

But, what happened with the 149 km he reaches before he retraced his path to the 177 km marker?*

OpenStudy (shane_b):

It doesn't count in displacement. You can think of displacement as the direct-line distance between two points. In the end, his displacement from the 257KM marker was only 80KM...it doesn't matter what path he took to get there. If the question asked for distance, it would be a different story.

OpenStudy (anonymous):

When I enter the answer 80 km (since it is an online homework system), the system says that it is incorrect.